Solve the system;
3x+y=15
4x+y=19

Answer by checkley71(8403)

3X+Y=15
-4X-Y=-19 THIS EQUATION IS MULTIPLIED BY -1 TO ELIMINATE THE Y TERM.
----------------------
-X=-4
X=4 ANSWER.
3*4+y=15
12+Y=15
Y=15-12
Y=3 ANSWER
-4*4+3=19
19=19